OpenStudy (anonymous):

If A = {1,2} and B = {3,4} , write all the relations from A to B.

OpenStudy (anonymous):

No of relation = 2^(a*b) = 2^4 = 16

OpenStudy (anonymous):

u should have 16 relations!

OpenStudy (anonymous):

Just find them

OpenStudy (anonymous):

i know that but am only getting 14 relations.

OpenStudy (anonymous):

u included \[\phi \]

OpenStudy (anonymous):

yea i did

OpenStudy (anonymous):

just show ur work we will find out where u missed

OpenStudy (anonymous):

okay ..

OpenStudy (anonymous):

A = {1, 2} and B = {3, 4} ∴A × B = {(1, 3), (1, 4), (2, 3), (2, 4)} ⇒ n(A × B) = 4 We know that if C is a set with n(C) = m, then n[P(C)] = 2m. Therefore, the set A × B has 24 = 16 subsets. These are Φ, {(1, 3)}, {(1, 4)}, {(2, 3)}, {(2, 4)}, {(1, 3), (1, 4)}, {(1, 3), (2, 3)}, {(1, 3), (2, 4)}, {(1, 4), (2, 3)}, {(1, 4), (2, 4)}, {(2, 3), (2, 4)}, {(1, 3), (1, 4), (2, 3)}, {(1, 3), (1, 4), (2, 4)}, {(1, 3), (2, 3), (2, 4)}, {(1, 4), (2, 3), (2, 4)}, {(1, 3), (1, 4), (2, 3), (2, 4)}

OpenStudy (anonymous):

okay yu finishd b4 me .. :P

OpenStudy (anonymous):

Did u understand!

OpenStudy (anonymous):

yea .. am checking where i went wrong

OpenStudy (anonymous):

u probably should have missed something

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

{(1, 3), (2, 3)},{(1, 3), (2, 4)} i missedd those .. tnx for the help :)

OpenStudy (anonymous):

welcome....))